OpenStudy (anonymous):

how do I solve the polynomial equation x^3+5x=9x

OpenStudy (anonymous):

First you want to always set a polynomial equation equal to 0, so the best move first is to subtract 9x from both sides: x^3+5x=9x turns into x^3-4x=0 Now we can forget about the 0 for a bit and focus on x^3-4x. Now these both share one solitary x so we can simply factor this out giving us: x(x^2-4) At this point it is solved as a difference of squares. Would you like for me to explain this too?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would that turn into x(x-2) (x-2) ?

OpenStudy (anonymous):

A difference of squares will always have opposite signs between the binomials, so flip one of the negative signs and you have it!

OpenStudy (anonymous):

If you want to, you can always check yourself by multiplying out of the FOIL method the O and I. If these two terms cancel each other out, the difference of squares was achieved correctly.

OpenStudy (anonymous):

so, then it would be x(x+2) (x-2) ?

OpenStudy (anonymous):

Oh, okay. Thanks

OpenStudy (anonymous):

Yes it would

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help with a different equation too? x(x+8)=x^2+6x+9 ? I know that first you have to equal it to zero, but after I equal it to zero I am left with 2x+9 and I don't know what to do next.

OpenStudy (anonymous):

@flixoe